# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v2

@@ 15 Appendix 3: Visualization and Accessibility of Oriented Graphs
In exploring oriented graphs, it became evident that fully grasping these structures often requires more than static illustrations in a paper can offer. While oriented graphs can be defined and studied abstractly, their complexity and behavior are best understood through dynamic, interactive representations. The Seymour Second Neighborhood Conjecture, in particular, necessitated defining paths, locating squares, and identifying degree-doubling nodes—tasks that involved intricate arc counting and relationships that benefit from visualization.
This realization motivated the creation of a complementary website. The site serves as a tool to enhance the accessibility of key definitions, lemmas, and proofs by providing visualizations and interactive examples. Our goal is to bridge the gap between abstract theory and intuitive understanding, making the intricate aspects of the conjecture approachable for a wider audience.
For example, consider the concept of a degree-doubling node or a lemma describing staircase-like structures in oriented graphs. While the paper describes these ideas in formal mathematical terms, the website illustrates them dynamically. This approach allows readers to ”see” the behavior and implications of these structures, offering a visual intuition that complements the formal reasoning presented in the paper.
This methodology draws inspiration from James Robert Brown’s insights in Philosophy of Mathematics, where he emphasizes the dual role of visualization in mathematics—as both a powerful tool for intuition and a potential pitfall if over-relied upon. By carefully balancing these aspects, we aim to enhance comprehension without compromising the rigor essential to mathematical proofs.
The website also provides an opportunity to include numerous examples and extended visual explanations that are not feasible within the space constraints of a traditional journal format. Through this medium, we invite readers to engage more interactively with the conjecture and its underlying theories, fostering a deeper understanding and encouraging constructive feedback.
Built using technologies like D3.js, JavaScript, and PHP, the website features dynamic visualizations of proof steps, interactive diagrams, and supplementary explanations. By demonstrating the synergy between mathematical results and computational tools, we aim to make these concepts not only accessible but also engaging for a broader audience.
@@
While the website is already live and functional, it is an evolving project. Additional examples, applications, and refinements are being added continuously.
Readers are invited to explore the website and are encouraged to provide feedback on any aspects of the site or the content itself, including suggestions for improvements, corrections, or additional examples.
The goal is to create a resource that not only supports this paper but also serves as a tool for further exploration and discussion in the field of oriented graph theory, the Seymour Second Neighborhood Conjecture, and the Graph Level Order (GLOVER) data structure. Thank you for your engagement and contributions to this ongoing effort.

### v3

@@ 15 Appendix 3: Visualization and Accessibility of Oriented Graphs
When we take a course on graph theory, oriented graphs are probably not on the course load. Second neighbors or squares of graphs are probably not either. To fully grasp this problem though we need to be able to delve into these definitions through visualizations.
For example, research on the Seymour Second Neighborhood Conjecture led this paper into the terms decreasing sequence property, induced subgraphs, containers, interior neighbors, exterior neighbors, and back arcs. The paper addressed how these terms were intertwined with one another. One example is where the paper says that there is a monotonically decreasing sequence.
James Robert Brown’s book Philosophy of Mathematics speaks of the importance of visualizations in mathematics. Proofs are still the foundation, but sometimes visualizations can serve as an inspiration or a spark for those proofs.
This realization motivated the creation of a complimentary website. The site serves as an interactive tool to understand research. It also has things like mathjax installed to read lemmas, definitions and proofs from this paper. Where possible, things like HTML canvas and JavaScript D3.js told are used to visualize these concepts and help improve understanding.
@@
This is an evolving project. Additional examples and applications will continuously be added
Readers are invited to explore the website. Please provide feedback on any aspects of the site or the content for improvements.
Our goal here was to help users connect with the Seymour Second Neighborhood Conjecture. Hopefully, this tool not only serves as support for this paper, but as an educational resource in the fields of graph theory and computer science. Thank you for your engagement